OpenStudy (anonymous):

The length of a triangle is increasing at a rate of 8 cm/s and its its width is increasing at a rate of 3 cm/s. When the length is 20 cm and the width is 10cm, how fast is the area of the rectangle increasing?

OpenStudy (dumbcow):

are we talking about a rectangle or triangle ??

OpenStudy (anonymous):

rectangle...

OpenStudy (dumbcow):

\[A = L*W\] find dA/dt using product rule \[\frac{dA}{dt} = \frac{dL}{dt}*W + L*\frac{dW}{dt}\]

OpenStudy (anonymous):

what numbers get plugged in where?

OpenStudy (dumbcow):

L is length , W is width dL/dt is rate length is increasing dW/dt is rate width is increasing